Exercise 1. Every rational number can be expressed as a quotient p/q so that p and q has no common factors.


1
Expert's answer
2021-01-13T18:28:59-0500

By definition, a rational number "x" is a quotient of two integers, we will note them as "a, b", where "b\neq 0". If "a=0", then "x=0 = \frac{0}{1}" is convenient. If "a\neq0," we note "r=gcm(a,b)\neq0", we have "x = \frac{a}{b}= \frac{a/r}{b/r}", where "a\/r, b\/r" are integers (as "r|a, r|b"), "b/r\neq 0" and "gcm(a\/r, b\/r)=1" (as if it was not the case, that would contradict to the fact that "r" is the greatest common divisor). Therefore "p=a\/r,q= b\/r" are integers that have no common factors and "x=\frac{a}{b}= \frac{a/r}{b/r}=\frac{p}{q}"